# Information Flow Routes: Automatically Interpreting Language Models at Scale

### v1

@@ Abstract
Information flows by routes inside the network via mechanisms implemented in the model. These routes can be represented as graphs where nodes correspond to token representations and edges to operations inside the network. We automatically build these graphs in a top-down manner, for each prediction leaving only the most important nodes and edges. In contrast to the existing workflows relying on activation patching, we do this through attribution: this allows us to efficiently uncover existing circuits with just a single forward pass. Additionally, the applicability of our method is far beyond patching: we do not need a human to carefully design prediction templates, and we can extract information flow routes for any prediction (not just the ones among the allowed templates). As a result, we can talk about model behavior in general, for specific types of predictions, or different domains. We experiment with Llama 2 and show that the role of some attention heads is overall important, e.g. previous token heads and subword merging heads. Next, we find similarities in Llama 2 behavior when handling tokens of the same part of speech. Finally, we show that some model components can be specialized on domains such as coding or multilingual texts.
@@
${}^{\ast}$ork done during an internship at Meta AI.
@@ 1 Introduction
Current state-of-the-art LMs are built on top of the Transformer architecture (Vaswani et al., 2017; Brown et al., 2020; Zhang et al., 2022; Chowdhery et al., 2022; Touvron et al., 2023a, b). Inside the model, each representation evolves from the current input token embedding [^1] Sometimes, along with positional encoding. to the final representation used to predict the next token. This evolution happens through additive updates coming from attention and feed-forward blocks. The resulting stack of same-token representations is usually referred to as “residual stream” (Elhage et al., 2021), and the overall computation inside the model can be viewed as a sequence of residual streams connected through layer blocks. Formally, we can see it as a graph where nodes correspond to token representations and edges correspond to operations inside the model (attention heads, feed-forward layers, etc.).
@@
While during model computation all the edges (i.e., model components) are present, computations important for each prediction are likely to form only a small portion of the original graph (Voita et al. (2019); Wang et al. (2023); Hanna et al. (2023), among others). We extract this important subgraph in a top-down manner by tracing information back through the network and, at each step, leaving only edges that were relevant (Figure 1). To understand which edges are important, we rely on an attribution method (Ferrando et al., 2022) and refuse from activation patching [^2] Originally introduced by Vig et al. (2020) to analyze LMs through the lens of causal mediation analysis. typical for the existing mechanistic interpretability workflows (Wang et al., 2023; Hanna et al., 2023; Conmy et al., 2023; Stolfo et al., 2023; Heimersheim & Janiak, 2023). Firstly, patching requires human efforts to create templates and contrastive examples and, therefore, is only applicable to a few pre-defined templates. Secondly, explaining a single prediction demands a substantial number of interventions (patches). Given that each intervention needs a forward pass, studying large models becomes increasingly impractical. In contrast, our method is about 100 times faster.
@@ 2 Extracting Information Flow Routes
Figure 3: General-case algorithm for extracting the important subgraph of an information flow graph.
@@
In this view, we can see a graph where (i) nodes correspond to token representations and (ii) edges correspond to operations inside the network that move information across nodes. Specifically, $x_{pos}^{lA}$ and $x_{pos}^{l}$ are representations of the token at position $pos$ after applying only attention block in layer $l$ or the entire layer, respectively. Each of $x_{1}^{l-1}$, …, $x_{pos}^{l-1}$ is connected to $x_{pos}^{lA}$ via attention edges (and via a residual stream edge from $x_{pos}^{l-1}$ to $x_{pos}^{lA}$), $x_{pos}^{lA}$ is connected to $x_{pos}^{l}$ via two edges: via feed-forward layers (FFNs) and via residual stream.
@@
While during model computation all the edges from Figure 2 are present, computations important for each prediction are likely to form only a small portion of the original graph (Voita et al. (2019); Wang et al. (2023); Hanna et al. (2023), etc). We extract this important subgraph in a top-down manner by tracing information back through the network (Figure 3). We start building the important subgraph from a single node – the representation on top of the residual stream. Then, we go over immediately connected lower nodes and, if they were important at this step (i.e., contributed sufficiently), add them to the graph along with the corresponding edges. Note that for $x_{pos}^{l}$, immediate connections contain $x_{pos}^{lA}$ with two different edges: via FFN and via residual connection, and we treat these connections separately.
@@
In the algorithm above, what is left is to define how to compute the importance of an edge. While lately it has become rather typical to use patching (Wang et al., 2023; Hanna et al., 2023; Conmy et al., 2023; Stolfo et al., 2023; Heimersheim & Janiak, 2023), instead of patching, we choose to use attribution. This choice is crucial for our work and makes our method about 100 times faster than alternatives while (i) being able to recover previously discovered circuits, (ii) doing this in a more versatile manner, and (iii) leading to many new observations.
In the rest of this section, we explain the specific attribution method we use.
@@ 2.1 Evaluating Edge Importance
For the attribution method, we adopt ideas from ALTI (Aggregation of Layer-Wise Token-to-Token Interactions) by Ferrando et al. (2022). While originally this method propagated attributions from output to input tokens through the entire model, we take only its part that defines contributions between immediately connected nodes. We choose this method because it is straightforward, easy to implement, and was shown effective for practical applications, e.g. detecting hallucinations in neural machine translation (Dale et al., 2023a, b; Guerreiro et al., 2023).
@@ 2.1.1 Importance in General Case
Note that in our graph (Figure 2), each node can be represented as a sum of the vectors corresponding to the incoming edges. According to ALTI, importance of one of the vectors (edges) to the overall sum (node) is proportional to its proximity to the resulting sum. Formally, if
${\bm{x}}={\bm{y}}_{1}+\dots+{\bm{y}}_{m},\vspace{-2ex}$
then
$\!\!\!\!importance({\bm{y}}_{j},{\bm{x}})\!=\!\frac{proximity({\bm{y}}_{j},{\bm{x}})}{\sum\limits_{k}proximity({\bm{y}}_{k},{\bm{x}})},$ (1)
$proximity({\bm{y}}_{j},{\bm{x}})=\max(-||{\bm{y}}_{j}-{\bm{x}}||_{1}+||{\bm{x}}||_{1},0).$
Here, we use negative distance as a measure of similarity: the smaller the distance between ${\bm{y}}_{i}$ and ${\bm{x}}$, the higher the impact of ${\bm{y}}_{i}$ on ${\bm{x}}$. Note also that we ignore contributions of the vectors lying beyond the $l_{1}$ length of ${\bm{x}}$. For more details, see Ferrando et al. (2022).
@@ 2.1.2 Defining FFN Edges
For the FFN blocks, edge vectors are straightforward. In the notation of Figure 2,
${\bm{x}}_{pos}^{l}$ $={\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\bm{x}}_{pos}^{lA}}+{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\text{FFN}_{l}({\bm{x}}_{pos}^{lA})},$
and the two terms correspond to the edges via residual connection and via FFN, respectively.
@@ 2.1.3 Defining Attention Edges
Attention edges correspond to connections between residual streams. To get vector updates coming from other residual streams to the current node, we follow previous work and decompose the output of an attention block into sum of vectors corresponding to its inputs (Kobayashi et al., 2020; Ferrando et al., 2022). We illustrate this decomposition in Figure 4.
Individual attention heads.
@@
$Attn^{h}(\mathbf{x}_{\leq i})=\sum_{j\leq i}\alpha_{i,j}^{h}f^{h}({\bm{x}}_{j}),$ (2)
where $f^{h}({\bm{x}}_{j})={\bm{x}}_{j}LW^{h}_{OV}$, $W_{OV}^{h}=W_{V}^{h}W_{O}^{h}$ the values and output combined matrix for head $h$, $\alpha_{i,j}^{h}$ are scalar attention weights, and $L$ is the linearized layer normalization (see Appendix A). [^3] From here onwards, we omit layer indices for readability.
In a bit more detail, a typical attention implementation (1) weights corresponding value vectors within each attention head, (2) concatenates head outputs, and (3) further multiplies by the output matrix $W_{O}$. In our view above, we split the output matrix $W_{O}$ into its head-specific parts $W_{O}^{h}$, drag these parts inside attention heads, and combine them with the head’s values matrix: $W_{OV}^{h}=W_{V}^{h}W_{O}^{h}$.
Overall, Figure 4 shows that on the way from attention head input to its output, each representation ${\bm{x}}_{j}$ is transformed linearly into ${f^{h}({\bm{x}}_{j})={\bm{x}}_{j}LW_{OV}^{h}}$ and multiplied by a scalar attention weight $\alpha_{i,j}^{h}$. This gives us the “raw output” emitted by each input vector ${\bm{x}}_{j}$ when treating attention weights as prediction-specific constants. In this view, information flows through attention heads by independent channels $\alpha_{i,j}^{h}f^{h}({\bm{x}}_{j})$ that converge in the next residual stream state.
@@
Entire attention block.
Figure 5 shows information flow for the entire attention block. Similarly to individual attention heads, here information also flows via independent channels corresponding to input tokens:
$Attn(\mathbf{x}_{\leq i})=\sum\limits_{h}^{H}\sum_{j\leq i}\alpha_{i,j}^{h}f^{h}({\bm{x}}_{j}).$ (3)
Following the general view from Section 2.1.1, we compute the importance of each of the terms in this sum individually and then aggregate edge capacities connecting the same pair of nodes. That is, sum over heads within each token, and additionally include a residual connection for the current token (Figure 5). Formally, the final attention edge importances are as follows:
${e_{i,j}=\left\{\begin{array}[]{@{}ll@{}}\sum_{h}^{H}e^{h}_{i,j}&\mbox{if}~{}j\neq i\\ \sum_{h}^{H}e^{h}_{i,j}+e^{r}_{i}&\mbox{if}~{}j=i\end{array}\right.}$
where $e_{i,j}^{h}$ are token-head importances and $e_{i}^{r}$ is the residual connection importance. Formally, ${e^{h}_{i,j}=importance(\alpha_{i,j}^{h}f^{h}({\bm{x}}_{j}),{\bm{x}}^{A}_{i})}$ and ${e^{r}_{i}=importance({\bm{x}}_{i},{\bm{x}}^{A}_{i})}$ respectively, as defined in equation (1).
@@ 2.2 Extracting the Important Subgraph
In the previous Section 2.1 we explained how we compute the importances of the edges in the full information flow graph (Figure 2) that are later used in Algorithm 3 to extract the important subgraph. The algorithm requires setting a threshold $\tau$ for the minimum edge importance in the graph which will complete the algorithm. One can make this algorithm more fine-grained by taking into account a few considerations.
Minimal head importance requirement.
First, we can exclude attention edges composed only of unimportant heads. Intuitively, if between two nodes none of the attention heads contributed significantly, then this connection might be considered unimportant. In this case, in equation (4) we set to zero the connections with edge values below $\tau$ and then renormalize the rest. Either of the two versions of the algorithm (with zeroing-out unimportant sub-edges or not) can be used – in our experiments, we did not notice significant differences. In the main text, we show results with this strategy.
Different thresholds for edge types.
Generally, feed-forward blocks have higher importance than attention heads. This is expected: while attention heads are only a (small) part of the attention block, the feed-forward block is not decomposed further. Therefore, one can set different thresholds for retaining attention and FFN edges. However, we do not do this in our experiments.
@@ 3 Information Flow vs Patching Circuits
First, we compare the information routes extracted with our method to the circuits found in previous work for two tasks: Indirect Object Identification (IOI) (Wang et al., 2023) and Greater-than tasks (Hanna et al., 2023). The IOI task consists of predicting the next word in sentences like “When Mary and John went to the store, John gave a drink to ___”. An initial clause features the names of two individuals (Mary and John). Then, the second clause depicts a person exchanging an item with the other person. The goal is to predict the second name in the main clause, i.e. the Indirect Object (Mary). In the Greater-than task, the model is prompted to predict a number given a sentence following the template: “The [noun] lasted from the year XXYY to the year XX___”. Here, the task is to predict a number higher than YY.
@@
The circuits for these tasks were previously found using activation patching and relying on “baseline” (contrastive) templates. These templates are designed to bring to the surface the specific task. For example, for IOI a contrastive template contains three different names instead of repeating one of them (Figure 7). One of the differences between information flow and patching results is that our method finds all the components that contributed to the prediction, while patching is claimed to find what is important for the original task but not the contrastive baseline. As we will see, (i) our method can also be used in this manner, and (ii) it gives more reasonable results.
@@ 3.1 Indirect Object Identification
For the IOI task, the previously found circuit contains several attention heads with various functions, e.g. Name Mover Heads, Duplicate Token Heads, etc. (Wang et al., 2023).
Using our method, we extract information flow routes for both the original task and the contrastive templates, and evaluate the activation frequency of attention heads. When looking at overall contributions, Figure 7 (left) shows that our routes largely consist of the heads discovered previously. There are, however, several heads which are always part of the routes (have near 1 activation frequency) but are not in the “patching circuit” (shown with pale circles). Interestingly, when looking at the difference with contrastive templates (Figure 7, right), these generic heads disappear. This makes sense: information flow routes contain all the components that were important for prediction, and these contain (i) generic components that are overall important, and (ii) task-specific heads. Note also how our method goes further than patching for identifying the difference between the original and the contrastive tasks. Figure 7 shows that e.g. the previous token heads (yellow) are important for both types of predictions and are not specific to the IOI task.

### v2

@@ Abstract
Information flows by routes inside the network via mechanisms implemented in the model. These routes can be represented as graphs where nodes correspond to token representations and edges to computations. We automatically build these graphs in a top-down manner, for each prediction leaving only the most important nodes and edges. In contrast to the existing workflows relying on activation patching, we do this through attribution: this allows us to efficiently uncover existing circuits with just a single forward pass. Unlike with patching, we do not need a human to carefully design prediction templates, and we can extract information flow routes for any prediction (not just the ones among the allowed templates). As a result, we can analyze model behavior in general, for specific types of predictions, or different domains. We experiment with Llama 2 and show that the role of some attention heads is overall important, e.g. previous token heads and subword merging heads. Next, we find similarities in Llama 2 behavior when handling tokens of the same part of speech. Finally, we show that some model components can be specialized on domains such as coding or multilingual texts. [^1] The proposed method is implemented in: https://github.com/facebookresearch/llm-transparency-tool.
@@
Work done during an internship at Meta AI.
@@ 1 Introduction
Current state-of-the-art LMs are built on top of the Transformer architecture (Vaswani et al., 2017; Brown et al., 2020; Zhang et al., 2022; Chowdhery et al., 2022; Touvron et al., 2023a, b). Inside the model, each representation evolves from the current input token embedding to the final representation used to predict the next token. This evolution happens through additive updates coming from attention and feed-forward blocks. The resulting stack of same-token representations is usually referred to as “residual stream” (Elhage et al., 2021), and the overall computation inside the model can be viewed as a sequence of residual streams connected through layer blocks. Formally, we can see it as a graph where nodes correspond to token representations and edges to operations inside the model (attention heads, feed-forward layers, etc.).
@@
While during a forward pass all the edges are present, computations important for each prediction are likely to form a small portion of the original graph (Voita et al., 2019; Wang et al., 2023; Hanna et al., 2023, among others). We extract this important subgraph in a top-down manner by tracing information back through the network and, at each step, leaving only edges that were relevant (Figure 1). To understand which edges are important, we rely on an attribution method (Ferrando et al., 2022) and refuse from activation patching [^2] Originally introduced by Vig et al. (2020) to analyze LMs through the lens of causal mediation analysis., typical for the existing mechanistic interpretability workflows (Wang et al., 2023; Hanna et al., 2023; Conmy et al., 2023; Stolfo et al., 2023). Firstly, patching requires human efforts to create templates and contrastive examples, thus is only applicable to a few pre-defined templates. Secondly, explaining a single prediction demands a substantial number of interventions (patches). Given that each intervention needs a forward pass, studying large models becomes increasingly impractical. In contrast, our method is about 100 times faster.
@@ 2 Extracting Information Flow Routes
Figure 3: General-case algorithm for extracting the important subgraph, the information flow routes, from the full information graph (Figure 2).
@@
In the graph (i) nodes correspond to token representations and (ii) edges to operations inside the network moving information across nodes. Specifically, $x_{pos}^{lA}$ and $x_{pos}^{l}$ are representations of the token at position $pos$ after the attention block in layer $l$ or the entire layer, respectively. Each $x_{1}^{l-1}$, …, $x_{pos}^{l-1}$ is connected to $x_{pos}^{lA}$ via attention edges (and via a residual stream edge from $x_{pos}^{l-1}$ to $x_{pos}^{lA}$), $x_{pos}^{lA}$ is connected to $x_{pos}^{l}$ via two edges: the FFN output and the residual stream.
@@
While during a forward pass all edges in Figure 2 are present, computations relevant for each prediction are likely to form a small portion of the original graph (Voita et al., 2019; Wang et al., 2023; Hanna et al., 2023, etc). We extract this important subgraph, the information flow routes, in a top-down manner by tracing information back through the network (Figure 3). We start from a single node – the representation on top of the residual stream. Then, we go over immediately connected lower nodes and, if important at this step, we add them to the subgraph along with the corresponding edges. The algorithm requires setting a threshold $\tau$ for the minimum edge importance.
@@
To complete the proposed algorithm (Figure 3), we need to specify how to compute the importance of an edge. While lately it has become typical to use patching (see Appendix A for a more formal description) (Wang et al., 2023; Hanna et al., 2023; Conmy et al., 2023; Heimersheim & Janiak, 2023), instead of patching, we choose to use attribution. This choice is crucial for our work and makes our method about 100 times faster than alternatives while (i) being able to recover previously discovered circuits, (ii) doing this in a more versatile manner, and (iii) leading to new observations. Next, we explain the specific attribution method we use.
@@ 2.1 Evaluating Edge Importance
For the attribution method, we adopt ideas from ALTI (Aggregation of Layer-Wise Token-to-Token Interactions) (Ferrando et al., 2022). While ALTI propagates attributions throughout the entire model, we only use its definition of contributions between connected nodes. We choose this method due to its simplicity, ease of implementation, and demonstrated effectiveness in practical applications, e.g. detecting hallucinations in neural machine translation (Dale et al., 2023a, b; Guerreiro et al., 2023).
@@ 2.1.1 Importance in General Case
In our graph (Figure 2), each node represents a sum of incoming vectors (edges). According to ALTI, the importance of each vector (edge) to the overall sum (node) is proportional to its proximity to the resulting sum. Formally, if ${\bm{y}}={\bm{z}}_{1}+\dots+{\bm{z}}_{m}$,
$\!\!\!\!importance({\bm{z}}_{j},{\bm{y}})\!=\!\frac{proximity({\bm{z}}_{j},{\bm{y}})}{\sum\limits_{k}proximity({\bm{z}}_{k},{\bm{y}})},$ (1)
$proximity({\bm{z}}_{j},{\bm{y}})=\max(-||{\bm{z}}_{j}-{\bm{y}}||_{1}+||{\bm{y}}||_{1},0).$
Here, we use negative distance as a measure of similarity: the smaller the distance between ${\bm{z}}_{j}$ and ${\bm{y}}$, the more the information of ${\bm{z}}_{j}$ in ${\bm{y}}$. Note also that we ignore contributions of the vectors lying beyond the $l_{1}$ length of ${\bm{y}}$. For more details, see Ferrando et al. (2022).
@@ 2.1.2 Defining FFN Edges
For the FFN blocks, edge vectors are straightforward. Following the notation of Figure 2:
${\bm{x}}_{pos}^{l}$ $={\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\bm{x}}_{pos}^{lA}}+{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\text{FFN}_{l}({\bm{x}}_{pos}^{lA})},$
where the terms correspond to the edges of the residual connection and FFN, respectively.
@@ 2.1.3 Defining Attention Edges
We follow previous work (Kobayashi et al., 2020; Ferrando et al., 2022) and decompose the output of an attention block into a sum of vectors (edges), each corresponding to a connection between residual streams (Figure 4).
Attention heads.
@@
$Attn^{h}(\mathbf{x}_{\leq pos})=\sum_{j\leq pos}\alpha_{pos,j}^{h}f^{h}({\bm{x}}_{j}),$ (2)
where $f^{h}({\bm{x}}_{j})={\bm{x}}_{j}LW^{h}_{OV}$, $W_{OV}^{h}=W_{V}^{h}W_{O}^{h}$ the values and output combined matrix for head $h$, $\alpha_{pos,j}^{h}$ are scalar attention weights, and $L$ is the linearized layer normalization (see Appendix B.1). [^3] From here onwards, we omit layer indices for readability.
Expanding further, a typical attention implementation (1) weights corresponding value vectors within each attention head, (2) concatenates head outputs, and (3) further multiplies by the output matrix $W_{O}$. We split the output matrix $W_{O}$ into its head-specific parts $W_{O}^{h}$, drag these parts inside attention heads, and combine them with the head’s values matrix: $W_{OV}^{h}=W_{V}^{h}W_{O}^{h}$.
Overall, Figure 4 shows that along an attention head each representation ${\bm{x}}_{j}$ is transformed linearly into ${f^{h}({\bm{x}}_{j})={\bm{x}}_{j}LW_{OV}^{h}}$ and multiplied by a scalar attention weight $\alpha_{pos,j}^{h}$. This gives us the “raw output” emitted by each input vector ${\bm{x}}_{j}$ when treating attention weights as prediction-specific constants. In this view, information flows through attention heads by independent channels $\alpha_{pos,j}^{h}f^{h}({\bm{x}}_{j})$ that converge in the next residual stream state, we refer to each of this channels as a sub-edge.
@@
Attention block.
The information in an attention block flows through all the independent channels (sub-edges) in the $H$ heads (Figure 5):
$Attn(\mathbf{x}_{\leq pos})=\sum\limits_{h}^{H}\sum_{j\leq pos}\alpha_{pos,j}^{h}f^{h}({\bm{x}}_{j}).$ (3)
We compute the importance of each of the sub-edges in this sum as described in Section 2.1.1 and aggregate across heads the capacities of those sub-edges connecting the same pair of nodes, $\sum_{h}^{H}e^{h}_{pos,j}$. Additionally, we include the importance of the residual connection for the current token, $e_{pos}^{\text{res\_attn}}$ (Figure 5). Formally, attention edge importances are computed as
${e^{\text{attn}}_{pos,j}=\left\{\begin{array}[]{@{}ll@{}}\sum_{h}^{H}e^{h}_{pos,j}&\mbox{if}~{}j\neq pos\\ \sum_{h}^{H}e^{h}_{pos,j}+e_{pos}^{\text{res\_attn}}&\mbox{if}~{}j=pos\end{array}\right.}$
where ${e^{h}_{pos,j}=importance(\alpha_{pos,j}^{h}f^{h}({\bm{x}}_{j}),{\bm{x}}^{A}_{pos})}$ and ${e_{pos}^{\text{res\_attn}}=importance({\bm{x}}_{pos},{\bm{x}}^{A}_{pos})}$ respectively, as defined in equation (1).
@@ 2.2 Extracting the Important Subgraph
In Section 2.1 we explained how we compute the importances of the edges in the full information flow graph (Figure 2). Finally, when building the important information flow subgraph (routes), we add only the edges with an importance above the specified threshold $\tau$ (Figure 3). Although we didn’t notice significant differences, in our experiments we first remove the sub-edges with importances $e^{h}_{pos,j}$ and $e^{\text{res\_attn}}_{pos,j}$ below $\tau$, and renormalize the rest before aggregating across heads.
Generally, feed-forward blocks have higher importance than attention heads. This is expected: while attention heads are only a (small) part of the attention block, the feed-forward block is not decomposed further. Therefore, one can set different thresholds for retaining attention and FFN edges, although we did not experiment with this.
@@ 3 Information Flow vs Patching Circuits
First, we compare our information routes to the circuits found in previous work for the Indirect Object Identification (IOI) (Wang et al., 2023) and Greater-than (Hanna et al., 2023) tasks. The IOI task consists of predicting the next word in sentences like “When Mary and John went to the store, John gave a drink to ___”. An initial clause features the names of two individuals (Mary and John). Then, the second clause depicts a person exchanging an item with the other person. The goal is to predict the second name in the main clause, i.e. the Indirect Object (Mary). In the Greater-than task, the model is prompted to predict a number given a sentence following the template: “The [noun] lasted from the year XXYY to the year XX___”. Here, the task is to predict a number higher than YY.
@@
The circuits for these tasks were previously found using activation patching with contrastive templates. These templates are designed to bring to the surface the specific task. For example, for IOI a contrastive template contains three different names instead of repeating one of them (Figure 7). One of the differences between information flow and patching results is that our method finds all the components that contributed to the prediction, while patching finds what is important for the original task but not the contrastive baseline. As we will see, (i) our method can also be used in this manner, and (ii) it gives more reasonable results.
@@ 3.1 Indirect Object Identification
For the IOI task, the previously found circuit contains several attention heads with various functions, e.g. Name Mover Heads, Duplicate Token Heads, etc. (Wang et al., 2023). Using our method, we extract information flow routes for both the original task and the contrastive templates, and evaluate the activation frequency of attention heads [^4] We consider an attention head activated if it has at least a sub-edge in the information flow routes (important subgraph).. When looking at overall contributions, Figure 7 (left) shows that our routes largely consist of the heads discovered previously. There are, however, several heads which are always part of the routes (have near 1 activation frequency) but are not in the “patching circuit” (shown with pale circles). Interestingly, when looking at the difference with contrastive templates (Figure 7, right), these generic heads disappear. This makes sense: information flow routes contain all the components that were important for prediction, and these contain (i) generic components that are overall important, and (ii) task-specific heads. Note also how our method goes further than patching for identifying the difference between the original and the contrastive tasks. Figure 7 shows that e.g. the previous token heads (yellow) are important for both types of predictions and are not specific to the IOI task.